# Vibrationally Assisted Exciton Transfer in Open Quantum Systems with Long-Range Interactions

### v1

@@ VII Discussion and Outlook
In this work, we characterized the role of entanglement and delocalization in long-range interacting spin chains interacting with an engineered reservoir. We found that the spin-spin interaction defines a specific class of entangled states that optimize the transport of excitations in the presence of an engineered environment. This conclusion still holds when accounting for static disorder and thermal excitations.
Our work is motivated by very recent experimental advances [36, 37, 46, 44, 84] and aims at providing a blueprint for the analog quantum simulation of Frenkel-exciton models based on trapped-ion hardware. Our work is a step towards the direct trapped-ion simulation of vibrationally assisted exciton transfer in chemical and biological complexes where the role of quantum coherence and delocalization in electron transfer is still debated [24, 32, 22, 25, 26, 27, 28, 29]. A trapped-ion quantum simulator offers, on one hand, full tunability of both the coherent and dissipative part of the evolution while, on the other hand, granting access to site-resolved observables that are crucial to understanding the energy flow dynamics [22].
@@
In this work, we focused on a spin chain coupled to a single, damped phonon mode. However, electron transfer in realistic systems often involves multiple vibrational modes. Therefore, the inclusion of multiple damped modes of motion will be paramount for the simulation of quantum dynamics of complex excitonic models. This can be achieved natively on trapped-ion quantum hardware, where multiple coolant ions can be used to control the individual cooling rates and the temperature of multiple bosonic modes. This capability supports the engineering of structured spectral densities [57], paving the way for studying non-Markovian effects in quantum dynamics [45, 5, 4]. Furthermore, the realization of Frenkel-exciton models with multiple modes with engineered spectral density will allow simulation across diverse vibrational frequency ranges, capturing both inter- and intra-molecular vibrations and low-frequency solvent environments [32, 59].
Including more than one bosonic mode in this model will also allow the simulation of multiply bridged donor-acceptor molecules [86], where intermediate sites experience distinct environments or couplings to different vibrational modes. Additionally, this setup provides an opportunity to introduce tunable anharmonic couplings among bosonic modes in the trapped-ion chain [87], facilitating the study of anharmonic effects, which are believed to influence electron transfer [88].
Here, we explored the effect of static disorder, but in realistic systems, electronic and motional dephasing are also present and play a key role [63]. In trapped-ion systems, electronic dephasing can be mimicked by introducing noisy light shifts [89, 78], and motional dephasing can be realized by controlling the trap frequencies [90]. While our study focuses on excitation-conserving $\sim\sigma_{z}(a^{\dagger}+a)$ spin-phonon coupling, real systems feature relaxation of excitations, which limits the transfer efficiency. In trapped-ion systems, it is possible to include spin relaxation experimentally by using a red-sideband (Jaynes-Cummings type) drive with dissipation.
@@ Appendix B Couplings
$J_{TT}=\left(\frac{1}{2}\frac{J}{(d+1)^{p}}+\frac{J}{(2+d)^{p}}+\frac{1}{2}\frac{J}{(3+d)^{p}}\right),$
@@
Where the distance between neighboring ions is set equal to 1 and $d$ is the number of coolants. The ratio of the different couplings in Eq. (B) for different values of $p$ and $d$ is shown in Fig. 7.
Figure 7: Inter-monomer couplings in the triplet-singlet basis: (a) Ratio of the different inter-dimer couplings as a function of $p$ for $d=2$. The horizontal gray dashed line is used as a guide to highlight the point where $J_{TT}^{2}/(J/3^{p})^{2}=1$ at $p\approx 3$. (b) Ratio of the different inter-dimer couplings as a function of $d$ for $p=1$.
@@
For $p=1$, results in Fig. 7(b) show that the ratio between $J_{TT}$ and any other coupling becomes larger as $d$ is increased. This means that increasing $d$ results in a more dominant triplet-triplet transfer, of course, at the cost of the transfer becoming slower. This trend is independent of the value of $p$. In the main text, the chosen values of $p=1$ and $d=2$ ensure that the triplet-triplet transfer is the most dominant while still having a fast transfer.
Appendix C Configurations with static disorder
@@
Appendix D Higher complexity monomers and more monomers
@@
$|\mathcal{E}_{D}^{1}\rangle$ is the state with the largest overlap with the $W$ state $|W_{D}\rangle$ and also the state with the highest eigenenergy. Similarly, equivalent results for states $|\mathcal{E}^{1,2,3}_{A}\rangle$ can be found for the acceptor sites in terms of states $|4\rangle$, $|5\rangle$, and $|6\rangle$. The eigenenergies of these states are identical to those in Eq. (D) but with negative on-site energy $\epsilon\rightarrow-\epsilon$. $|\mathcal{E}_{A}^{1}\rangle$ also corresponds to the acceptor state with the highest energy and largest overlap with $|W_{A}\rangle$.
@@
Appendix E Experimental realization of the vibrational and vibronic coupling terms
Here we ouline in detail the derivation of the trapped-ion Hamiltonian corresponding to the vibrational and vibronic coupling terms in Eq. (1) via the proposed experimental scheme in Sec. VI.1 and the discussion on the effect of the unwanted interactions arising from this scheme.

### v2

@@ VII Discussion and Outlook
In this work, we characterized the role of entanglement and delocalization in long-range interacting spin chains interacting with an engineered reservoir. We found that the spin-spin interaction defines a specific class of entangled states that optimize the transport of excitations in the presence of an engineered environment. This conclusion still holds when accounting for static disorder, noise, and thermal excitations.
Our work is motivated by very recent experimental advances [36, 37, 46, 44, 90] and aims at providing a blueprint for the analog quantum simulation of Frenkel exciton models based on trapped-ion hardware. Our work is a step towards the direct trapped-ion simulation of vibrationally assisted exciton transfer in chemical and biological complexes, where the role of quantum coherence and delocalization in electron transfer is still debated [24, 32, 22, 25, 26, 27, 28, 29]. A trapped-ion quantum simulator offers, on the one hand, full tunability of both the coherent and dissipative parts of the evolution while, on the other hand, granting access to site-resolved observables that are crucial to understanding the energy flow dynamics [22].
@@
In this work, we focused on a spin chain coupled to a single, damped phonon mode. However, electron transfer in realistic systems often involves multiple vibrational modes. Therefore, the inclusion of multiple damped modes of motion will be paramount for the simulation of quantum dynamics of complex excitonic models. This can be achieved natively on trapped-ion quantum hardware, where multiple coolant ions can be used to control the individual cooling rates and the temperature of multiple bosonic modes. This capability supports the engineering of structured spectral densities [60], paving the way for studying non-Markovian effects in quantum dynamics [45, 5, 4] where strong, ultra-strong, and time-dependent system-environment couplings can be studied. Furthermore, the realization of Frenkel exciton models with multiple modes with engineered spectral density will allow simulation across diverse vibrational frequency ranges, capturing both inter- and intra-molecular vibrations and low-frequency solvent environments [32, 64].
Including more than one bosonic mode in this model will also allow the simulation of multiply bridged donor-acceptor molecules [98], where intermediate sites experience distinct environments or couplings to different vibrational modes [99, 100]. Additionally, this setup provides an opportunity to introduce tunable anharmonic couplings among bosonic modes in the trapped-ion chain [101], facilitating the study of anharmonic effects, which are believed to influence electron transfer [102].
In the work presented here, we focused on the role that coherent superpositions of states within each monomer play in determining the transfer efficiency. Considering more than two monomers, as we did in Section V, opens the possibility of studying the transfer of an excitation initially delocalized over multiple monomers. The presence of different types of quantum coherences and their survival over time has been found to be related to the high efficiency of excitation transfer in photosynthetic complexes [24, 103]. Previous studies [104, 105, 106, 31] have pointed out that the consideration of non-Markovian baths is essential to model correctly the survival of these coherences, which, as pointed out earlier, represents an interesting future research direction both experimentally and theoretically.
Here, we explored the effect of static disorder and white-noise fluctuations of the energy shifts through electronic dephasing, the latter as well as motional dephasing are common noise sources in realistic setups and play a key role in transfer dynamics [69, 72, 107]. In trapped-ion systems, electronic dephasing can be mimicked by introducing noisy light shifts [70, 91], and motional dephasing can be realized by controlling the trap frequencies [108]. As shown in Ref. [70], more complicated noise profiles can be realized in these platforms, opening the possibility to study richer coherence dynamics. While our study focuses on excitation-conserving $\sim\sigma_{z}(a^{\dagger}+a)$ spin-phonon coupling, real systems feature relaxation of excitations, which limits the transfer efficiency. In trapped-ion systems, it is possible to include spin relaxation experimentally by using a red-sideband (Jaynes-Cummings type) drive with dissipation.
@@ Appendix B Couplings
$J_{TT}=\left(\frac{1}{2}\frac{J}{(1+d)^{p}}+\frac{J}{(2+d)^{p}}+\frac{1}{2}\frac{J}{(3+d)^{p}}\right),$
@@
Where the distance between neighboring ions is set equal to 1, and $d$ is the number of coolants. The ratio of the different couplings in Eq. (B) for different values of $p$ and $d$ is shown in Fig. 7.
Figure 7: Inter-monomer couplings in the triplet-singlet basis: (a) Ratio of the different inter-dimer couplings as a function of $p$ for $d=2$. The horizontal gray dashed line is used as a guide to highlight the point where $J_{TT}^{2}/(J/3^{p})^{2}=1$ at $p\approx 3$. (b) Ratio of the different inter-dimer couplings as a function of $d$ for $p=1$. (c) Ratio of the different inter-dimer couplings as a function of $d$ for $p=3$.
@@
For $p=1$, the results in Fig. 7(b) show that the ratio between $J_{TT}$ and any other coupling becomes larger as $d$ is increased. This means that increasing $d$ results in a more dominant triplet-triplet transfer, of course, at the cost of the transfer becoming slower. This trend is independent of the value of $p$. For example, the case $p=3$ representing the dipole interaction, which is native in many systems, shows a similar trend as a function of $d$, as shown in Fig. 7(c). In the main text, the chosen values of $p=1$ and $d=2$ ensure that the triplet-triplet transfer is the most dominant while still having a fast transfer.
Appendix C Numerical simulation details
Since the dynamics described in the excitation transfer involve a non-zero number of excitations in the phonon mode, especially in the case of higher temperatures, the numerical integration of Eq. (2) requires choosing an appropriate cutoff number of phonon excitations $N_{\rm ph}$. In Fig. 8, we show how the evolution of different observables is affected by the choice of this cutoff.
Figure 8: Phonon cutoff dependence: Time evolution of observables for different values of the phonon cutoff $N_{\rm ph}$, represented by distinct curves. The chosen parameters are the same as the parameter set $p_{1}$ in Table. 2. (a) Population in the donor as a function of time. (b) Population in the phonon mode as a function of time. The black dashed line represents $\alpha^{2}=(g/2\omega)^{2}=0.25$, which is the expected value of the population in the steady state.
As depicted in the figure, for these parameters, a cutoff of $N_{\rm ph}>10$ excitations already provides an appropriate size of the phonon mode Hilbert space, as no significant changes can be noticed in the observables when the value is increased to $N_{\rm ph}=15$ (the two curves lie on top of each other). We adjust the phonon cutoff accordingly for the different parameter sets used in this work.
Appendix D Entanglement transfer
One of the main results in this work is that the initialization of the system in an entangled (delocalized) state of the qubits in the donor site speeds up the transfer into the acceptor site. Under ideal conditions (i.e., at low temperature and on resonance), the steady state in the acceptor is also an entangled state, as illustrated in Fig. 2. For example, this is evidenced by the fact that $P_{D}\approx 1-P_{T_{A}}$ when the system is initialized in the electronic entangled state $|T_{D}\rangle$. However, we make this argument more concrete here, where we explicitly show the dynamics of the qubit’s entanglement in terms of the concurrence.
For a two-qubit state, pure or mixed, the concurrence can be used to determine the entanglement. Mathematically, the concurrence is defined as [115, 116]:
$\mathcal{C}_{AB}(\rho)=\text{max}(0,\mu_{1}-\mu_{2}-\mu_{3}-\mu_{4})\,,$ (25)
where $\rho$ is the quantum state of the two qubits A and B, and $\mu_{j}$ are the eigenvalues of the matrix $\tilde{\rho}=\rho S\rho^{*}S$, with $S=\sigma_{y}\otimes\sigma_{y}$. Here, $\sigma_{y}$ is the conventional Pauli matrix in the y direction. The eigenvalues are defined in descending order $\mu_{j}>\mu_{j+1}$. If $\mathcal{C}_{AB}(\rho)=0$, the state has no entanglement (separable state), and if $\mathcal{C}_{AB}(\rho)=1$, the system is maximally entangled. In Fig. 9, we show the evolution of the concurrence between the donor qubits $\mathcal{C}_{12}$ as well as the concurrence between the acceptor qubits $\mathcal{C}_{34}$ for the two monomer case, described in section III.
Figure 9: Entanglement transfer: Time evolution of the concurrence between different qubits and the purity of the state. (a) The chosen parameters are the ones in parameter set $p_{1}$ in Table. 2. (b) Same as in (a) but with $\bar{n}=0.1$. (c) Same as in (a) but with $\epsilon=0.26\omega$.
[…]
In the ideal conditions of panel (a), the state of the phonons is the displaced vacuum state, so the steady state of the total system is the pure state $|\psi_{T_{D},0}\rangle$. Consequently, $\rho^{2}\rightarrow 1$ at long times, as reported in the plot. If we increase $\bar{n}$, as in panel (b), the steady state can be a mixed state of multiple phonon states, as shown by $\rho^{2}<1$. Nonetheless, the electronic state steady state can still be purely $|T_{D}\rangle$, and in that case, all the initial entanglement between qubits in the donor site is still transferred to the acceptor site.
In panel (c), we move away from resonance by setting $\epsilon=0.26\omega$, such that the initial state $|T_{D}\rangle$ can weakly couple to $|S_{A}\rangle$. In that case, the electronic steady state is a mixture of $|T_{A}\rangle$ and $|S_{A}\rangle$, consequently, not all the initial entanglement is transferred to the acceptor, namely, $\mathcal{C}_{34}<1$ in the steady state.
When the donor and acceptor sites contain more than two qubits (see section V), other entanglement measures different from concurrence would be needed to characterize the entanglement dynamics of the mixed multi-qubit states. However, we expect similar entanglement dynamics. For instance, for the parameters chosen here, if the system is initialized in $|\mathcal{E}_{D}^{1}\rangle$, ideal conditions would lead to transfer into the state $|\mathcal{E}_{A}^{1}\rangle$ (see Fig. 5(b)). This state has a high overlap with state $|W_{A}\rangle$, as shown in Appendix F, which implies that the steady state of the system has a large entanglement. For instance, if we consider the donor/acceptor to have $N$ qubits, with $N$ being even, and consider the bipartition of $N/2$ qubits in each subsystem, the bipartite entanglement that can be transferred from the donor to the acceptor grows approximately as $S\propto\log_{2}(N)+C$ [81] when the system is initialized in $|\mathcal{E}_{D}^{1}\rangle$.
Appendix E Configurations with static disorder
@@
Appendix F Higher complexity monomers and more monomers
@@
$|\mathcal{E}_{D}^{1}\rangle$ is the state with the largest overlap with the $W$ state $|W_{D}\rangle$ and also the state with the highest eigenenergy. Similarly, equivalent results for states $|\mathcal{E}^{1,2,3}_{A}\rangle$ can be found for the acceptor sites in terms of states $|4\rangle$, $|5\rangle$, and $|6\rangle$. The eigenenergies of these states are identical to those in Eq. (F) but with negative on-site energy $\epsilon\rightarrow-\epsilon$. $|\mathcal{E}_{A}^{1}\rangle$ also corresponds to the acceptor state with the highest energy and largest overlap with $|W_{A}\rangle$.
@@
Appendix G Experimental realization of the vibrational and vibronic coupling terms
Here we outline in detail the derivation of the trapped-ion Hamiltonian corresponding to the vibrational and vibronic coupling terms in Eq. (1) via the proposed experimental scheme in Sec. VI.1 and the discussion on the effect of the unwanted interactions arising from this scheme.